OpenStudy (ammarah):

The area of a triangle is 338 square yards. The height of the triangle is four times its base. Find the height and base of the triangle.

OpenStudy (ammarah):

@Mertsj

OpenStudy (anonymous):

area=1/2bh 338=1/2(x)(4x) 676=4x^2 169=x^2 X=13, so the base is 13 and height is 52

OpenStudy (ammarah):

thankyou!!! can u help me with some more?

OpenStudy (anonymous):

yeah

OpenStudy (ammarah):

how did u get 676?

OpenStudy (anonymous):

Divide both sides by 1/2 to get rid of the 1/2 in the formula, its the same as multiplying by 2.

OpenStudy (ammarah):

Ohhkk

OpenStudy (anonymous):

yeeeaaahhh

OpenStudy (ammarah):

I need help 72cm cubed =______ m^